# 3D Gaussian Splatting: Survey, Technologies, Challenges, and Opportunities

### v1

@@ 5.3 Editable 3D Gaussian Splatting
3DGS, with its advantages of real-time rendering, representation of complex scenes, and explicit representation, has naturally attracted considerable attention from researchers focusing on 3DGS editing. Unfortunately, current editable 3DGS works often lack precise training supervision, which poses a significant challenge for editing. In this section, we categorize existing works according to different tasks.
5.3.1 Manipulation by Text
To address this challenge, the existing works can be classified into two distinct categories. The first type introduces the score distillation loss as in Eq. 3. Unlike AIGC 4.2, these methods require editing prompts as additional conditions to guide the editing process.
Based on the SDS, GaussianEditor [162] introduces semantic control into 3DGS editing, enabling semantic-based tracking and automatic masking of editing regions. More importantly, this work proposes a hierarchical 3DGS and multi-generation anchor loss, which stabilizes the editing process and mitigates the SDS’s impact of randomness. Additionally, GaussianEditor introduces 2D inpainting techniques to provide guidance for the tasks of object removal and incorporation. Following Dreamgaussian [75], GSEdit [163] uses the pre-trained Instruct-Pix2Pix [164] model instead of the image generation model to compute score distillation loss for 3DGS editing.
The second type focuses on sequentially editing multi-view 2D images before reconstructing 3DGS.
GaussianEditor2 [165] leverages multi-modal models, large language models, and segmentation models to predict the editable regions from given textual descriptions. It then optimizes the relevant Gaussian primitives within the target regions based on the images edited by the 2D editing models.
However, such a paradigm introduces an intuitive problem: how to ensure consistency in multi-view editing.
@@ 5.3.1 Manipulation by Text
GaussCtrl [166] introduces a depth-guided image editing network, ControlNet [167], utilizing its ability to perceive geometry and maintain multi-view consistency in the editing network. It also introduces a latent code alignment strategy in the attention layers, ensuring that the edited multi-view images remain consistent with the reference images.
The work [136] aims to introduce inverse rendering and rendering in 3D latent space to maintain consistency on the attention map. Furthermore, it introduces an editing consistency module and an iterative optimization strategy, further enhancing the multi-view consistency and editing capabilities.
Unlike editing methods for 3DGS, recent discussions have increasingly focused on editing 4DGS.
Recent work, Control4D [168], delves into this area by introducing 4D GaussianPlanes, which structurally decompose the 4D space to ensure consistency across spatial and temporal dimensions as Tensor4d. Based on the GaussianPlanes, it designs a 4D generator using super-resolution generative adversarial networks [169], which learns a generation space on GaussianPlanes from the edited images generated by the diffusion model and employs a multi-stage progressive guiding mechanism to enhance the local-global quality.
5.3.2 Manipulation by Other Conditions
In addition to text-controlled editing, existing works have explored 3DGS editing methods under various conditions. TIP-Editor [170] requires the provision of an editing text, a reference image, and the editing location to finely control 3DGS. The core technologies include a stepwise 2D personalization strategy for learning the existing scene and the new content separately, and a coarse-fine editing strategy for accurate appearance. This approach allows users to perform various editing tasks, such as object insertion and stylization.
Point’n Move [171] requires the user to provide annotated points for the objects to be edited. It achieves controlled editing (including inpainting the removal area) of the objects through dual-stage segmentation, inpainting, and recomposition steps.
5.3.3 Stylization
@@ 5.3.4 Animation
As described in 5.1, some dynamic 3DGS works, such as SC-GS [150], can achieve animation effects by animating sparse control points. AIGC-related works, such as BAGS [111], aim to utilize video input and generation models to animate existing 3DGS. Similar research has also been mentioned in the context of Human Reconstruction.
Additionally, CoGS [174] discusses how to control this animation. Based on dynamic representations [135, 143], it uses a small MLP to extract relevant control signals and align the deformation of each Gaussian primitive. And then, CoGS generates 3D masks for the regions to be edited to reduce unnecessary artifacts.
5.4 Semantic Understanding
Endowing 3DGS with semantic understanding capabilities allows for the extension of 2D semantic models into 3D space, thereby enhancing the model’s comprehension in 3D environments. This can be applied to various tasks such as 3D detection, segmentation, and editing.
Many works attempt to leverage pre-trained 2D semantic-aware models for extra supervision on semantic attributes.
Early work, Feature 3DGS [175], distills pre-trained 2D foundational models to jointly construct 3DGS and feature fields. By introducing a parallel feature rasterization strategy and regularization, it endows the 3DGS with the capability of spatial understanding and enables the design of Promptable Explicit Scene Representations for downstream tasks.
Subsequently, Gaussian Grouping [176] introduces the concept of Gaussian groups and extends Identity Encoding attributes to achieve it. This work proposes treating multi-view data as a video sequence with gradually changing views and utilizes a pre-trained object tracking model [177] to ensure multi-view consistency of the segmentation labels obtained from SAM (Segment Anything)[178]. Additionally, Gaussian groups are supervised in both 2D and 3D spaces and applied directly for editing. Similarly, the work[179] addresses semantic inaccuracy by introducing KNN clustering and Gaussian filtering, which can constrain nearby Gaussians and eliminate distant Gaussians.
CoSSegGaussians [180] leverages a pre-trained point cloud segmentation model in conjunction with a dual-stream feature fusion module. This module integrates unprojected 2D features derived from 2D encoders with 3D features from 3D encoders, following the prediction of Gaussian locations [15]. By employing decoders and semantic supervision, CoSSegGaussians effectively imbues Gaussian primitives with semantic information.
Others [138, 181, 182] focus on incorporating text-visual alignment features for open-world understanding.
@@ 5.4 Semantic Understanding
A significant challenge is the high dimensionality of CLIP features, which makes direct training and storage difficult compared to original Gaussian attributes. The work [138] introduces corresponding continuous semantic vectors into the 3DGS by extracting and discretizing dense features from CLIP [183] and DINO [184], which are used to predict semantic indices $m$ in Discrete Feature Space by MLPs as in VQ-VAE [185]. It also proposes incorporating an uncertainty attribute to describe the instability and frequent changes of Gaussian primitives and designing an adaptive spatial smoothing loss to deliberately decrease the spatial frequency of the embedded compact semantic features.
LangSplat [181] compresses scene-specific CLIP features using a trained autoencoder to reduce training memory requirements. To achieve this objective, LangSplat introduces hierarchical semantics—subparts, parts, and wholes—constructed using SAM [178], which addresses point ambiguity across multiple semantic levels and facilitates scene understanding in response to arbitrary text queries. Subsequently, FMGS [182] mitigates the issue of large CLIP feature dimensions by introducing multi-resolution hash encoders [9].
5.5 Physics Simulation
@@
Recent efforts aim to extend 3DGS to simulation tasks. Based on the "what you see is what you simulate" philosophy, PhysGaussian [139] reconstructs a static 3DGS as the discretization of the scene to be simulated, and then incorporates continuum mechanics theory along with the Material Point Method (MPM) [186] solver to endow 3DGS with physical properties. To stabilize rotation-based varying appearances and fill particles into the void internal region, PhysGaussian proposes an evolving orientation and internal filling strategy.
6 Technical Classification
@@
3DGS can generally be divided into the following stages, as illustrated in Fig. 3: initialization, attribute optimization, splatting, regularization, training strategy, adaptive control, and post-processing. Additionally, some concurrent works aim to incorporate supplementary information and representations, thereby enhancing the capability of 3DGS. These technical improvements not only enhance the rendering performance of the original 3DGS but also address specific tasks in derivative works. Consequently, this section delves into the technological advancements in 3DGS, with the aim of providing valuable insights for researchers in related fields.
6.1 Initialization
@@
Proper initialization has been shown to be crucial, as it directly affects the optimization process [187]. The initialization of 3DGS is typically performed using sparse points derived from Structure-from-Motion (SfM) or through random generation. However, these methods are often unreliable, especially under weak supervision signals, such as sparse view settings and generative tasks.
Combining pre-trained models is an optional approach.
@@
Pre-training a 3D model on a limited number of 3D samples and using it as an initialization prior is a viable strategy [51]. This approach can enhance the performance of initialization to some extent, although its effectiveness is contingent upon the data used. To address this limitation, pretrained 3D generative models [80, 81, 87] or monocular depth estimation models [100, 102] are also introduced for initialization purposes. Additionally, some works [80] propose the introduction of new perturbation points to achieve a more comprehensive geometric representation.
Improving initialization strategies is also important.
@@
Based on the analysis of the role of SfM in capturing low-frequency signals within the spectrum, Sparse-large-variance (SLV) initialization is designed to effectively focus on the low-frequency distribution identified by SfM [187].
@@
Utilizing other representations can also enhance initialization capabilities. By determining the Local Volumes from a coarse parametric point model, a small number of Gaussians are initialized within each volume, thereby avoiding excessive assumptions about the target [188]. Subsequently, an initialization strategy based on Visual Hull is proposed, enabling the acquisition of structural priors from four images [57].
Discussion: Precise initialization has the potential to benefit various derivative works of 3DGS by enhancing performance and ensuring a stable training process.
6.2 Attribute Expansion
The original attributes of 3DGS include position, scale, rotation, Spherical Harmonic (SH) coefficients, and an opacity value. Some works have extended these attributes to make them more suitable for downstream tasks. It can be categorized into improvements of existing attributes or the introduction of novel attributes, as shown in Fig. 6.
6.2.1 Improving Attributes
Certain attributes of Gaussian can be custom-tailored, thereby making 3DGS suitable for a wider range of tasks.
Scale: By collapsing the $z$-scale to zero and incorporating additional supervision on depth, normal, or shell maps, the works [160, 137, 156, 32, 73] aim to improve Gaussian primitives to make them flatter and more suitable for surface reconstruction, thereby mitigating inaccuracies in Gaussian geometric reconstruction. The $z$ direction can be approximated as the normal direction. Similarly, a scale constraint, which limits the ratio of the major axis length to the minor axis length [139, 126, 116], ensures that the Gaussian primitives remain spherical to mitigate the issue of unexpected plush artifacts caused by overly skinny kernels.
@@ 6.2.1 Improving Attributes
Gaussian: By introducing shape parameters, an attempt is made to replace the original Gaussians with a Generalized Exponential (GEF) mixture [27]. Traditional 3DGS can be viewed as a special case of the Generalized Exponential Family (GEF) mixture ($\beta=2$), enhancing the representational efficiency of Gaussians,
@@ 6.2.2 Additional Attributes
By adding new attributes and supervisions, the original representation capabilities of 3DGS can be augmented.
@@
Semantic Attributes: By introducing semantic attributes and corresponding supervision, works, such as [113, 114, 122, 130, 175, 131], are endowed with enhanced spatial semantic awareness, which is crucial for tasks such as SLAM and editing. After the splatting process used for semantic attributes, the 3DGS’s semantic attributes are supervised using 2D semantic segmentation maps. Additionally, improving methods for extracting semantic information [171] and introducing high-dimensional semantic-text features, such as CLIP and DINO features [181, 138, 182], have been employed to address a wider range of downstream tasks. Similar to semantic attributes, the Identity Encoding attributes can group 3DGS that belong to the same instance or stuff, which is more effective for multi-object scenes [176].
@@
Attribute Distributions: Learning a distributed attribute with reparameterization techniques instead of a fixed value is an effective approach to prevent local minima in 3DGS [84] and mitigate its reliance on Adaptive Control of 3DGS [47]. In addition to these works focusing on the distribution prediction of position attributes, the distribution of scale attributes has also been incorporated [84]. By sampling the predicted attribute distributions, Gaussian primitives for splatting can be obtained.
@@
Temporal Attributes: Replacing the original static attributes with temporal attributes is key to animating the 3DGS [110, 109, 113, 152]. For 4D attributes, including rotation, scale, and position, existing works render 3DGS on timestep $t$ by either taking time slices at specific time points [152] or decoupling the $t$ dimension from 4D attributes [109, 110]. Also, the introduction of 4D SH is crucial for time-varying color attributes. For this, the Fourier series is typically used as the adopted basis functions to endow SH with temporal capabilities in existing works [113, 110]. Note that due to involving different timesteps, such attributes often require video-based training.
@@
Displacement Attributes: Displacement attributes are indispensable as they describe the relationship between the final and initial positions of Gaussian primitives. These attributes can be classified based on their dependency on time. Time-independent displacement attributes are often used to correct coarse position attributes, which can be directly optimized in the same manner as other attributes [46, 63]. Time-dependent displacement attributes can describe the positional changes of static 3DGS, thereby achieving dynamic representations. This approach often involves introducing a small MLP to predict displacement based on timestep $t$ [145, 143, 144] and other control signals[174].
@@
Physical Attributes: Physical attributes encompass a broad range of potential properties that describe the objective physical laws governing Gaussian primitives, thereby endowing 3DGS with more realistic representation. For instance, shading-related attributes such as diffuse color, direct specular reflection, residual color, and anisotropic spherical Gaussian can be utilized for specular reconstruction [37, 39, 38]. Additionally, shadow scalar is introduced to express shadows [146], and velocity is employed to represent the transient information of Gaussian primitives, which is essential for describing dynamic properties [106]. These attributes are typically updated by considering the influence of physical properties at specific rendering positions [37, 146, 39] or by incorporating supplementary supervised information, such as flow maps [106].
@@
Discrete Attributes: Utilizing discrete attributes in place of continuous attributes is an effective method for compressing high-dimensional representations or representing complex motion. This is often achieved by storing the index values of the VQ codebook [18, 23, 19, 17] or the motion coefficient for motion basis [147] as the discrete attributes of Gaussian primitives. However, discrete attributes can lead to performance degradation; combining them with compressed continuous attributes may be a potential solution [138].
@@
Inferred Attributes: This type of attribute does not require optimization; they are inferred from other attributes and used for downstream tasks. The Parameter-Sensitivity Attribute reflects the impact of parameter changes on reconstruction performance. It is represented by the gradient of the parameter and is used to guide compression clustering [19]. The Pixel-Coverage attribute determines the relative size of Gaussian primitives at the current resolution. It is related to the horizontal or vertical size of the Gaussian primitives and is used to guide the Gaussian’s scale to meet the sampling requirements of multi-scale rendering [34].
@@
Weight Attributes: The Weight Attributes rely on structured representations, such as Local Volumes [188], Gaussian-kernel RBF [150], and SMPL [189], which determine the attributes of query points by calculating the weights of related points.
@@
Other Attributes: The Uncertainty Attributes can help maintain training stability by reducing the loss weight in areas of high uncertainty [138, 55]. And, the ORB-Features Attributes, extracted from image frames [123], play a crucial role in establishing 2D-to-2D and 2D-to-3D correspondences [119].
@@
Discussion: The modification of Gaussian attributes facilitates the execution of a wider range of downstream tasks, offering an efficient approach as it obviates the need for additional structural elements. Moreover, the integration of new Gaussian attributes with supplementary information constraints also has the potential to significantly enhance the representational efficacy of the original 3DGS. For instance, semantic attributes can, in certain scenarios, yield more precise object boundaries.
6.3 Splatting

### v2

@@ V-C Editable 3D Gaussian Splatting
3DGS, with its advantages of explicit representation, has naturally attracted considerable attention from researchers focusing on 3DGS editing. Unfortunately, current editable 3DGS works often lack precise training supervision, which poses a significant challenge for editing. In this section, we categorize existing works according to different tasks.
V-C1 Manipulation by Text
To address this challenge, the existing works can be classified into two distinct categories. The first type introduces the score distillation loss as in Eq. 8. These methods require editing prompts as additional conditions to guide the editing process [201, 202]. GaussianEditor [201] enables semantic-controlled 3DGS editing through SDS, utilizing hierarchical 3DGS and anchor loss for stability, while incorporating 2D inpainting guidance for object manipulation. Following Dreamgaussian, GSEdit [202] uses the pre-trained Instruct-Pix2Pix [203] model instead of the image generation model for SDS. However, such methods remain constrained by pre-trained diffusion models, particularly when handling complex editing prompts.
@@ V-C1 Manipulation by Text
The second type focuses on editing multi-view images before reconstructing. GaussianEditor2 [204] employs multi-modal, language, and segmentation models to locate editable regions from text inputs, then optimizes targeted Gaussians based on 2D-edited images. However, this paradigm introduces an intuitive problem: how to ensure consistency in multi-view editing [205, 206]. GaussCtrl [205] introduces a depth-guided image editing network, ControlNet [207], utilizing its ability to perceive geometry and maintain multi-view consistency in the editing network. It also introduces a latent code alignment strategy in the attention layers, ensuring that the edited multi-view images remain consistent with the references.
Unlike editing methods for 3DGS, recent discussions have increasingly focused on editing 4DGS. Recent work, Control4D [208] leverages 4D GaussianPlanes to structurally decompose four-dimensional space, ensuring spatiotemporal consistency through Tensor4D representation, while incorporating a super-resolution GAN-based 4D generator [209] that learns from diffusion-generated edited images. However, it remains challenged on non-rigid movements.
V-C2 Manipulation by Other Conditions
In addition to text-controlled editing, existing works have explored 3DGS editing methods under various conditions, such as mixed conditions [210] and points [211]. TIP-Editor [210] enables fine-grained 3DGS control through text, reference image, and location inputs, utilizing stepwise 2D personalization and coarse-fine editing strategies to support diverse tasks like object insertion and stylization. And Point’n Move [211] enables object-level editing through point annotations, utilizing a dual-stage process of segmentation, inpainting, and recomposition, which demonstrates improved control capability. Recent research [212] has introduced a training-free 3DGS splitting paradigm that achieves editing plane control by formulating it as a constrained minimization problem, preserving visual fidelity through moment conservation while mitigating Gaussian overflow via an analytically derived closed-form solution.
V-C3 Stylization
@@ V-C4 Animation
As described in Sec. V-A, some dynamic 3DGS works, such as SC-GS [184], can achieve animation effects by animating sparse control points. AIGC-related works, such as BAGS [144], aim to utilize video input and generation models to animate existing 3DGS. Similar research has also been mentioned in the context of Human Reconstruction. Additionally, CoGS [187] discusses how to control animation. Based on dynamic representations [169, 174], it uses a small MLP to extract relevant control signals and align the deformation of each Gaussian primitives.
V-D Relightable
Relightable 3DGS has also emerged as one of the recent challenges gaining significant attention. Decoupling texture and lighting represents a common approach in relighting tasks. In early work, Relightable 3D Gaussian [215] and GS-IR [216] represent scenes using points with normal, BRDF, and decomposed lighting attributes for relighting. However, these methods face challenges in handling reflective scenes. Therefore, follow-up work [34] introduce accurate normal estimation and residual color terms to effectively model view-dependent reflections and complex lighting interactions. To address limitations in complex materials like semi-transparent volumes and furs, OLAT-GS [217] decomposes observed color into the attenuated light intensity, received incident illumination and scattering value. And then, GS3 [218] combines spatial and angular Gaussians with a triple splatting process to model geometry and reflectance properties, incorporating neural networks for self-shadowing and global illumination. Despite their capability in handling challenging geometry and appearance, further exploration is needed for transparent materials and indirect lighting.
V-E Semantic Understanding
Endowing 3DGS with semantic understanding capabilities allows for the extension of 2D semantic models into 3D space, thereby enhancing the model’s comprehension in 3D environments. This can be applied to various tasks such as 3D detection, segmentation, and editing. Many works attempt to leverage pre-trained 2D semantic-aware models for extra supervision on semantic attributes [219, 220, 221, 222]. Feature 3DGS [220] leverages pre-trained 2D models to create joint 3DGS and feature fields, enabling spatial understanding through feature rasterization and regularization for downstream promptable tasks. However, these approaches remain constrained by multi-view consistency and open-world perception challenges. Subsequent research [223, 224] aims to incorporate contrastive learning losses as auxiliary supervision to achieve interactive 3D segmentation. To mitigate the dependence on 2D pre-trained models in the aforementioned works, the work [225] attempts to leverage reconstructed Featured 3DGS to fine-tune the 2D pre-trained encoder.
@@ V-E Semantic Understanding
Others focus on incorporating text-visual alignment features for open-world understanding [226, 227, 228]. A significant challenge is the high dimensionality of CLIP features, which makes direct training and storage difficult compared to original Gaussian attributes. The work [226] introduces corresponding continuous semantic vectors into the 3DGS by extracting and discretizing dense features from CLIP [229] and DINO [230], which are used to predict semantic indices $m$ in Discrete Feature Space by MLPs as in VQ-VAE [231]. Subsequently, FMGS [228] mitigates the issue of large CLIP feature dimensions by introducing multi-resolution hash encoders [171].
V-F Physics Simulation
@@
Recent efforts aim to extend 3DGS to simulation. Based on the "what you see is what you simulate" philosophy, PhysGaussian [191] reconstructs a static 3DGS as the discretization of the scene to be simulated, and then incorporates continuum mechanics theory along with the Material Point Method [232] solver to endow 3DGS with physical properties. Similar approaches have also been explored in PhysDreamer [233]. Nevertheless, these physics-based simulations typically entail significant computational overhead. VR-GS [234] design an efficient interactive simulation system for VR, offering users a rich platform for 3D content manipulation.
VI 3DGS Components and Improvements
@@
3DGS can generally be divided into the following stages, as illustrated in Fig. 2: initialization, attribute optimization, splatting, regularization, training strategy, and adaptive control. In follow-up studies, researchers have enhanced and refined the fundamental 3DGS components. These technical improvements not only enhance the rendering performance of the original 3DGS but also address specific tasks in derivative works. Consequently, this section delves into the technological improvements in 3DGS, with the aim of providing valuable insights for researchers in related fields.
VI-A Initialization
@@
Proper initialization has been shown to be crucial, as it directly affects the optimization process [235]. The initialization of 3DGS is typically performed using sparse points derived from Structure-from-Motion (SfM) or through random generation. However, these methods are often unreliable, especially under weak supervision signals.
@@
Combining pre-trained models is an optional approach. Pre-training a 3D model on a limited number of 3D samples and using it as an initialization prior is a viable strategy [52]. This approach can enhance the performance of initialization to some extent, although its effectiveness is contingent upon the data used. To address this limitation, pretrained 3D generative models [100, 101, 105] or monocular depth estimation models [129, 130, 62] are also introduced for initialization. Based on these, the work [100] also introduces some perturbation points to achieve a more comprehensive geometric representation.
@@
Improving initialization strategies is also important. Based on the analysis of the role of SfM in capturing low-frequency signals within the spectrum, Sparse-large-variance initialization is designed to effectively focus on the low-frequency distribution identified by SfM [235].
@@
Utilizing other representations can also enhance initialization capabilities. By determining the Local Volumes from a coarse model, a small number of Gaussians are initialized within each volume, thereby avoiding excessive assumptions [236]. Similarly, some initialization based on Visual Hull [64], Flame [78] or Mesh [80, 237, 217] are proposed, enabling the acquisition of structural geometric priors.
Discussion: The incorporation of pre-trained or structured information during the initialization of 3DGS is crucial, as high-fidelity initialization can mitigate training instabilities, particularly in under-determined scenarios. It is feasible to determine initialization protocols according to the available prior knowledge.
VI-B Attribute Expansion
The original attributes of 3DGS include the position, scale, rotation, Spherical Harmonic (SH) coefficients, and opacity value. Some works have extended these attributes to make them more suitable for downstream tasks. It can be categorized into improvements of existing attributes or the introduction of novel attributes, as shown in Fig. 4.
VI-B1 Improving Attributes
Certain attributes of vanilla Gaussian can be custom-tailored, thereby making 3DGS suitable for a wider range of tasks.
Scale: By collapsing the $z$-scale to zero and incorporating additional supervision on depth, normal, or shell maps, the works [197, 199, 192, 26, 92] aim to improve Gaussian primitives to make them flatter and more suitable for surface reconstruction, where $z$ direction can be approximated as the normal direction. Conversely, a scale constraint, which limits the ratio of the major axis length to the minor axis length [191, 238, 150], ensures that the Gaussian primitives remain spherical to mitigate the issue of unexpected plush artifacts caused by overly skinny kernels.
@@ VI-B1 Improving Attributes
Gaussian: By introducing shape parameters, an attempt is made to replace the original Gaussians with a Generalized Exponential Family (GEF) mixture [17]. Traditional 3DGS can be viewed as a special case of the GEF mixture ($\beta=2$), enhancing the representational efficiency of Gaussians,
@@ VI-B2 Additional Attributes
By adding new attributes and corresponding supervisions, the original representation capabilities of 3DGS can be augmented.
@@
Semantic Attributes: By introducing them and corresponding supervision, works such as [147, 148, 160, 161, 220, 162, 225, 224, 223] are endowed with enhanced spatial semantic awareness, which is crucial for tasks such as SLAM and editing. After the semantic attributes’ splatting, the 3DGS’s semantic attributes are supervised using 2D semantic segmentation maps. Additionally, methods to improve the extraction of semantic information [211] and introducing high-dimensional semantic-text, such as CLIP and DINO features [227, 226, 228], have been employed to address a wider range of downstream tasks. Similar to semantic attributes, the identity encoding attributes can group 3DGS that belong to the same instance or stuff [219], which is more effective for multi-object scenes.
@@
Attribute Distributions: Learning position distributions with reparameterization techniques instead of a fixed value is an effective approach to prevent local minima [103] and mitigate its reliance on Adaptive Control of 3DGS [46]. In addition to these works focusing on the distribution prediction of position attributes, the distribution of the scale has also been incorporated [103]. By sampling the predicted distributions, Gaussian primitives can be obtained.
@@
Temporal Attributes: Replacing the original static attributes with temporal attributes is key to animating the 3DGS [188, 140, 147, 189]. For 4D attributes, including rotation, scale, and position, existing works render 3DGS on timestep $t$ by either taking time slices [189] or decoupling the $t$ dimension from 4D attributes [140, 188]. Also, the introduction of 4D SH is crucial for time-varying color attributes. For this, the Fourier series is typically used as the adopted basis functions to endow SH with temporal capabilities [147, 188]. Note that due to involving different timesteps, these attributes often require video-based training. This regularization primarily aims to improve attributes in Gaussian primitives [238, 26, 191, 98, 80], as in Sec. VI-B.
@@
Displacement Attributes: They can describe the relationship between the final and initial attributes in Gaussian primitives and be classified based on their condition. Condition-independent displacement attributes are often used to refine coarse attributes, which can be directly optimized in the same manner as other attributes [45]. Condition-dependent displacement attributes can describe the changes of static 3DGS, thereby achieving dynamic representations and controllable representations. This approach often involves introducing a small MLP to predict displacement based on timestep $t$ [176, 174, 175], expression and other control signals[187, 41, 79, 83, 76, 85, 86, 84, 70].
@@
Physical Attributes: They encompass a broad range of properties describing the physical laws governing Gaussian primitives, thus endowing 3DGS with more realistic representation. For instance, shading-related attributes like diffuse color, direct specular reflection, residual color, shadow, and anisotropic spherical Gaussian can be used for specular reconstruction and relighting [34, 36, 35, 217, 218, 177]. Additionally, the velocity attributes can represent the transient information of Gaussian, essential for describing dynamic objects [178]. These attributes are typically optimized by considering the influence of physical laws at specific rendering positions [34, 177, 36] or by incorporating supplementary information, such as flow maps [178].
@@
Discrete Attributes: Utilizing discrete attributes in place of continuous ones is an effective method for compressing high-dimensional representations and representing complex motion. This is often achieved by storing the index values of the VQ codebook [11, 14, 12, 10, 120] or the motion coefficient for motion basis [182] as the discrete attributes in Gaussian primitives. However, discrete attributes inevitable lead to performance degradation; combining them with compressed continuous attributes may be a potential solution [226].
@@
Inferred Attributes: These attributes do not require optimization; they are inferred from other attributes. The Parameter-Sensitivity attributes reflects the impact of parameter changes on reconstruction performance and are represented by the gradient of the parameter, guiding compression clustering [12]. The Pixel-Coverage attributes determines the relative size of Gaussian primitives at the current resolution. It is related to the horizontal or vertical size of the Gaussian primitives and guides their scale to meet sampling requirements in multi-scale rendering [30].
@@
Weight Attributes: They rely on structured representations, such as Local Volumes [236], Gaussian-kernel RBF [184], Mesh [237], and SMPL [239], to determine the attributes of query points by calculating the weights of structured points.
@@
Other Attributes: The Uncertainty Attributes can help maintain training stability by reducing the loss weight in areas with high uncertainty [226, 58]. The ORB-Features Attributes, extracted from image frames [157], play a crucial role in establishing 2D-to-2D and 2D-to-3D correspondences [152].
@@
Discussion: The modification of Gaussian attributes facilitates the execution of a wider range of downstream tasks, offering an efficient approach as it obviates the need for additional structural elements. Moreover, the integration of new attributes with supplementary information constraints also has the potential to significantly enhance the representational efficacy of the original 3DGS. For instance, semantic attributes can, in certain scenarios, yield more precise object boundaries.
VI-C Splatting